On Monday, Harold picked up six donuts and three large coffees for the office staff. He paid $6.39. On Tuesday, Melinda picked up four donuts and six large coffees for the office staff. She paid $7.50. What is the cost of one donut? What is the cost of one large coffee?
Answer by addingup(3677)

6d + 3c = 6.39 multiply all times 2 and subtract from the next equation
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
12d + 6c = 12.78
-
4d + 6c = 7.50
-----------------
8d + 0c = 5.28
8d = 5.28
d = 0.66
~~~~~~~~~~~~~~~~~~~~~~~~~~~~
6d + 3c = 6.39
6(0.66) + 3c = 6.39
3.96 + 3c = 6.39
3c = 2.43
c = 0.81
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Note: I used the first equation to solve for c, but you can use either one of the equations. Here's using the 2nd equation:
4d + 6c = 7.50
4(0.66) + 6c = 7.50
2.64 + 6c = 7.50
6c = 4.86
c = 0.81